# DecoR: Deconfounding Time Series with Robust Regression

### v1

@@ Abstract
Causal inference on time series data is a challenging problem, especially in the presence of unobserved confounders. This work focuses on estimating the causal effect between two time series, which are confounded by a third, unobserved time series. Assuming spectral sparsity of the confounder, we show how in the frequency domain this problem can be framed as an adversarial outlier problem. We introduce Deconfounding by Robust regression (DecoR), a novel approach that estimates the causal effect using robust linear regression in the frequency domain. Considering two different robust regression techniques, we first improve existing bounds on the estimation error for such techniques. Crucially, our results do not require distributional assumptions on the covariates. We can therefore use them in time series settings. Applying these results to DecoR, we prove, under suitable assumptions, upper bounds for the estimation error of DecoR that imply consistency. We show DecoR’s effectiveness through experiments on synthetic data. Our experiments furthermore suggest that our method is robust with respect to model misspecification.
@@ 1 Introduction and Related Work
In this paper, we assume that the confounder is sparse in the frequency domain (or, more generally, after a suitable basis transformation). This assumption allows us to frame the problem as an adversarial outlier problem in the frequency domain and thereby enabling us to apply robust regression techniques to estimate the causal effect reliably. Figure 1 illustrates our proposed method, Deconfounding by Robust regression (DecoR), and offers a graphical comparison with ordinary least squares (OLS). We analyze DecoR theoretically and provide assumptions under which DecoR is consistent for estimating the causal effect.
Our approach was inspired by Mahecha et al. (2010) who predict temperature sensitivity of ecosystem respiratory processes in the case where basal respiration, the unobserved confounder, is slowly varying. While Mahecha et al. (2010) also consider the time series data in the frequency domain, they assume that the support of the confounder is known (which is not required in our framework). Furthermore, they focus on the application and neither provide a formal mathematical framework nor any statistical guarantees. The work by Ćevid et al. (2020) and Scheidegger et al. (2023) also consider sparsity with hidden confounders. However, they consider i.i.d. data and assume a sparse parameter vector rather than sparse confounding. Consequently, their procedure is different from ours in that they apply singular value decomposition to the data and then trim the large eigenvalues. A related line of research focuses on spatial deconfounding in environmental and epidemiological applications, see e.g. Clayton et al. (1993); Paciorek (2010); Page et al. (2017). These works assume that a spatial regression problem is confounded by an unobserved confounder that is slowly varying. Different methods have been developed for solving this problem In Reich et al. (2006), Hughes and Haran (2013) and Prates et al. (2019) the residual spatial process is restricted to be orthogonal to the covariates. Another approach is to remove the slowly varying components from either the response or the covariates or both (Paciorek, 2010; Thaden and Kneib, 2018; Keller and Szpiro, 2020; Dupont et al., 2022). A similar methodology, that also includes removing the slowly varying part of the covariate, was developed by Sippel et al. (2019) for meteorological time series data. More recently, Guan et al. (2023) and Marques et al. (2022) consider Gaussian random fields as data generation process and propose to remove confounding by considering different scales when estimating the (unconfounded) covariance matrix.
@@
The remainder of this paper is structured as follows. We formally introduce the problem of causal inference with unobserved confounders for time series in Section 2. In Section 2.1 we present DecoR and showcase how it is used for causal inference in the presence of unobserved confounders. We provide theoretical guarantees for DecoR in Section 3.
@@ 2 Sparse Representation for Deconfounding
Let $d\in\mathbb{N}$ and let $T\in\mathbb{R}_{>0}$ denote a fixed time horizon. Let $X=(X_{t})_{t\in[0,T]}$ be a stochastic process in $\mathbb{R}^{d}$ and $U=(U_{t})_{t\in[0,T]}$ a stochastic process in $\mathbb{R}$. Let $\eta=(\eta_{t})_{t\in[0,T]}$ be a process of i.i.d. centered Gaussian random variables independent of $X$ and with constant variance $\sigma^{2}_{\eta}\geq 0$. Let $\beta\in\mathbb{R}^{d}$ and let $Y=(Y_{t})_{t\in[0,T]}$ be a stochastic process that satisfies
@@
Consider a (known) orthonormal basis $\phi=\{\phi_{k}\}_{k\in\mathbb{N}}$ of $L^{2}([0,T])$. We typically choose the cosine basis (see Definition C.1) as $\phi$ in our applications. The main assumption of this paper is that the confounder is sparse in this basis.
Definition 2.1 ($(\phi,G)$-sparse process ).
@@
The core idea is to consider the pairs $\{(T_{k}^{\phi,n}(X),T_{k}^{\phi,n}(Y))\}_{k\leq n}$ as a new data set and to observe that by Assumption 1 only the data points with an index $k\in G$ are confounded. If $G$ is known or if there exists a known subset $S\subseteq\mathbb{N}$ with $G\subseteq S$, the unconfounded dataset $\{(T_{k}^{\phi,n}(X),T_{k}^{\phi,n}(Y))\}_{k\notin S}$ can be analyzed using standard linear inference methods to, for example, consistently estimate $\beta$. However, in many cases, $G$ (or $S$) might be unknown. In the next section, we introduce DecoR, an algorithm for estimating $\beta$ without assuming knowledge of $G$ but assuming sparsity instead.
@@ 3 Theoretical Guarantees for Robust Regression
This work discusses two algorithms for estimating $\beta$ in (4), both of these methods include some form of linear regression. Algorithm 2 takes as input a collection $\mathcal{U}_{n}$ of candidate sets for the inliers $G_{n}^{\mathsf{c}}$. For example, we might have knowledge of an upper bound $a$ for the number of outliers. We can then define $\mathcal{U}_{n}=\{S\subseteq\{1,\dots,n\}\ |\ |S|\geq n-a\}$. We can then search over all elements in $\mathcal{U}_{n}$, compute the OLS estimate on the corresponding data and choose the estimate with the smallest prediction error. We call this method Brute Force Search (BFS) and detail it in Algorithm 2. We will see in Section 3.1 that if, indeed, $G_{n}^{\mathsf{c}}\in\mathcal{U}_{n}$, the procedure is consistent under suitable assumptions.
@@
while $\left\lVert e\right\rVert_{2}<\mathrm{err}$ do
@@
     $e\leftarrow|Y-X\hat{\beta}^{t}_{\texttt{Tor}}|$
     $S_{t}\leftarrow\mathrm{HT}(e,a)$triangleright see (6)
@@ 3.1 Guarantees for BFS
where $K>0$ is the constant from Lemma D.1, and define
@@
The proof can be found in Appendix D.2. Theorem 3.1 implies that BFS is consistent if the noise variance $\sigma^{2}$ is equal to 00 and $\min_{S\in\mathcal{U}_{n}}||X_{S\setminus G_{n}}||_{2}\neq 0$, implying that $\mathcal{U}_{n}$ does not contain any subsets of $G_{n}$. Furthermore, it shows that if the fraction of outliers is vanishing with increasing $n$, BFS is consistent under mild assumptions on the collection $\mathcal{U}_{n}$ of candidate sets; it suffices, for example, if the distribution of $X$ is such that, with high probability [^6] See Appendix C.4 for a formal definition.,
$\min_{S\in\mathcal{U}_{n}}||X_{S\setminus G_{n}}||_{2}\in\Omega(\sqrt{n}),\qquad\max_{S\in\mathcal{U}_{n}}||X_{S}||_{2}\in\mathcal{O}(\sqrt{n})$ (7)
@@
For details, see Appendix E.2. Conditions (8) and (7) hold, for example, when we have a known sequence $\{c_{n}\}_{n\in\mathbb{N}}$ such that, for all $n$, $|G_{n}|\leq c_{n}$ and $c_{n}$ $\in o(\sqrt{n}/\log(n))$, we define the collection of candidate sets as $\mathcal{U}_{n}\coloneqq\{S\in\mathcal{P}(\{1,\dots,n\})\mid|S|=n-c_{n}\}$ and the rows of $X$ are i.i.d. Gaussian, see Bhatia et al. (2015, Theorem 15). If the fraction of outliers is non-vanishing (and $c_{n}<n/2$), Theorem 3.1 implies that for a distribution of $X$ satisfying (7) and the same choice of candidate sets defined by $\mathcal{U}_{n}\coloneqq\{S\in\mathcal{P}(\{1,\dots,n\})\mid|S|=n-c_{n}\}$, that $|\hat{\beta}^{n}_{\texttt{BFS}}(\mathcal{U}_{n})-\beta|\in\mathcal{O}(\sigma)$. Theorem 3.1 implies uniform consistency results, too. For example, in the case of $\sigma=0$, we obtain uniform consistency even over classes of distributions with outliers that shift the original data points by an arbitrary amount. On the contrary, regression that is based on a robust loss function, such as the Huber loss (Huber, 1964), does not come with similar guarantees, not even with respect to pointwise consistency or in the case $\sigma=0$.
@@ 3.2 Guarantees for Torrent
In practice, we observe that Torrent requires only very few interrations until it convergences (cf. Table 8). We can use Lemma 3.2 to obtain an improved guarantee for Algorithm 3.
Theorem 3.3.
@@
$\eta\coloneqq\max_{S\subseteq\{1,\dots,n\}\text{ s.t. }|S|=a_{n}}\frac{\left\lVert X_{V(S)}\right\rVert_{2}}{\sqrt{\lambda_{\mathrm{min}}\left(X_{S}^{T}X_{S}\right)}}<\frac{1}{\sqrt{2}}.$
@@
$\left\lVert\hat{\beta}^{n,a_{n}}_{\texttt{Tor}}-\beta\right\rVert_{2}\leq\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\epsilon_{S_{t}}\right\rVert_{2}+\frac{2\left\lVert X_{V(S_{t})}\right\rVert_{2}\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\epsilon_{S_{t}}\right\rVert_{2}+\sqrt{2}\|\epsilon_{V(S_{t})}\|_{2}}{\sqrt{\lambda_{\mathrm{min}}\left(X_{S_{t}}^{T}X_{S_{t}}\right)}\left(1-\sqrt{2}\eta\right)}.$
@@
Corollary 3.4 (sub-Gaussian noise).
Assume the setting of Theorem 3.3 with i.i.d. zero-mean sub-Gaussian noise $\epsilon$ with variance proxy $\sigma^{2}$ and assume $c_{n}<n/2$. Then, there exists a constant $K>0$ such that for all $\delta>0$ with probability at least $1-\delta$
@@
While Corollary 3.4 states consistency under appropriate conditions and vanishing fractions of outliers, the results in Bhatia et al. (2015) are not sufficiently strong to imply consistency in this setting. Furthermore, even for non-vanishing fraction of outliers our bounds are in general tighter since the maximum over $S$ in the definition of $\eta$ is taken jointly over the denominator and numerator, not separately as in Bhatia et al. (2015).
@@ 4 Guarantees for DecoR
For the remainder of this section we assume Setting 1 and that Assumption 1 is satisfied for $G$ and $\phi$. More precisely, let $G\subseteq\mathbb{N}$, let $\phi=\{\phi_{k}\}_{k\in\mathbb{N}}$ be an orthonormal basis of $L^{2}([0,T])$, let $U=(U_{t})_{t\in[0,T]}\in\mathcal{M}^{2}_{[0,T]}$ be a $(\phi,G)$-sparse process and let $X=(X_{t})_{t\in[0,T]}$ be a stochastic process in $\mathbb{R}^{d}$. Let $\eta=(\eta_{j})_{j\in[0,T]}$ be a stochastic process of i.i.d. centered Gaussian random variables with variance $\sigma_{\eta}^{2}\geq 0$ and let $Y_{t}$ be a stochastic process, such that, for all $t\in[0,T]$,
@@
We also assume some regularity conditions on $\phi$ and $U$. Specifically, we assume that for all $k\in\mathbb{N}$ it holds that $\phi_{k}$ is right-continuous (or left-continuous) and that the trajectories of $U$ satisfy the same condition almost surely. This implies that, almost surely888 Without assuming right-continuity (9) only holds almost surely, almost everywhere; because we need to discretize the process we require equality almost surely for all $t\in[0,T]$. , for all $t\in[0,T]$
@@
    For all $n\in\mathbb{N}$ and $l,k\leq n$ it holds that $\frac{1}{n}\sum_{j=1}^{n}\phi_{l}(Tj/n)\phi_{k}(Tj/n)=\mathbbm{1}{\{l=k\}}$.
@@
    For every $\delta>0$ there exists $c^{\prime}>0$ and $\bar{n}\in\mathbb{N}$ such that for all $n\geq\bar{n}$ there exists $S^{\prime}_{n}\subseteq\{1,\dots,n\}$ with $|S^{\prime}_{n}|=2c_{n}+d$ such that for all $S^{\prime\prime}\subset S^{\prime}_{n}$ with $|S^{\prime\prime}|=d$ it holds that with probability at least $1-\delta$
@@
Assumption 2 (i) states that the chosen orthonormal basis maintains its orthogonality and normalization properties when applied to discretized observations; this is satisfied, for example, by the cosine basis or the Haar basis (Haar, 1909), see Appendix C. Together with (9), this implies (2). Assumption 2 (ii) requires that $X$ is non-sparse in the $\phi$-domain. This condition intuitively says that there is more information in $X$ than is lost by confounding. This condition is relatively mild and accommodates a wide array of commonly used stochastic processes, including Ornstein-Uhlenbeck processes, Brownian motions, and band-limited processes. We are now able to state the main theoretical results, the convergence properties of DecoR-BFS (DecoR with BFS as robust regression algorithm) and DecoR-Tor (DecoR with Tor as robust regression algorithm).
Theorem 4.1 (Convergence properties of DecoR-BFS).
Let $c_{n}$ be a known sequence of natural numbers such that $|G_{n}|\leq c_{n}$ and let Assumption 2 be satisfied. Assume that $d=1$, that $U$ is independent of $\epsilon$, that $\sup_{t\in[0,T]}\mathbb{E}[X_{t}^{2}]<\infty$ and that for all $n\in\mathbb{N}$ and $j,i\leq T$ it holds that $\frac{1}{n}\sum_{l=1}^{n}\phi_{l}(Tj/n)\phi_{l}(Ti/n)=\mathbbm{1}{\{i=j\}}$. Define a sequence of candidate sets by $\mathcal{U}_{n}:=\{S\subseteq\{1,\dots,n\}\ |\ |S|=n-c_{n}\}$. If DecoR is executed with BFS and the sequence of sample sets $\mathcal{U}_{n}$, then there exists $\bar{n}\in\mathbb{N}$ such that for all $n\geq\bar{n}$ with high probability
$\left|\hat{\beta}^{\phi,n}_{\texttt{DecoR}}-\beta\right|\in\mathcal{O}\left(\sigma_{\eta}\left(\frac{c_{n}\log(n/c_{n})}{n}\right)^{1/4}\right).$
Many commonly used bases, such as the cosine basis and the Haar basis, satisfy the conditions of Theorem 4.1.
Theorem 4.2 (Convergence properties of DecoR-Tor).
@@
If DecoR is executed with Torrent and the sequence of threshold parameters $a_{n}=n-c_{n}$, then there exists $\bar{n}\in\mathbb{N}$ such that for all $n\geq\bar{n}$ with high probability
$\left\lVert\hat{\beta}^{\phi,n}_{\texttt{DecoR}}-\beta\right\rVert_{2}$ $\in\mathcal{O}\left(\sigma_{\eta}\sqrt{\frac{c_{n}\log(n/c_{n})}{n}}\right).$
If the assumptions of Theorem 4.1 or Theorem 4.2 are satisfied and either there is no noise, that is, $\sigma_{\eta}=0$, or the number of confounded components $c_{n}$ grows more slowly than the sample size $n$, that is, $c_{n}\in o(n)$, then DecoR produces a consistent estimator for $\beta$. Conversely, if there is non-zero noise and the number of confounded components $c_{n}$ is asymptotically proportional to $n$, that is, $c_{n}\sim n$, the estimation error of DecoR is asymptotically bounded by the noise variance $\sigma_{\eta}$, up to constant factors. While DecoR-Tor has polynomial-time complexity and in our experiments converges very fast (cf. Table 8), the theoretical result (Theorem 4.2) contains stronger theoretical assumptions than the one based on BFS (Theorem 4.1) , specifically it requires (10) to hold. In contrast, DecoR-BFS does not require (10), but its computational complexity is exponential in $n$ when $c_{n}$ is not bounded by a constant.

### v2

@@ Abstract
Causal inference on time series data is a challenging problem, especially in the presence of unobserved confounders. This work focuses on estimating the causal effect between two time series that are confounded by a third, unobserved time series. Assuming spectral sparsity of the confounder, we show how in the frequency domain this problem can be framed as an adversarial outlier problem. We introduce Deconfounding by Robust regression (DecoR), a novel approach that estimates the causal effect using robust linear regression in the frequency domain. Considering two different robust regression techniques, we first improve existing bounds on the estimation error for such techniques. Crucially, our results do not require distributional assumptions on the covariates. We can therefore use them in time series settings. Applying these results to DecoR, we prove, under suitable assumptions, upper bounds for the estimation error of DecoR that imply consistency. We demonstrate DecoR’s effectiveness through experiments on both synthetic and real-world data from Earth system science. The simulation experiments furthermore suggest that DecoR is robust with respect to model misspecification.
KEYWORDS:
Causal inference, confounding, Earth system science, robust regression, time series analysis
@@ 1 Introduction and Related Work
In this paper, we assume that the confounder is sparse in the frequency domain (or, more generally, after a suitable basis transformation). This assumption allows us to frame the problem as an adversarial outlier problem in the frequency domain and thereby enabling us to apply robust regression techniques to estimate the causal effect reliably. Figure 1 illustrates our proposed method, Deconfounding by Robust regression (DecoR), and offers a graphical comparison with ordinary least squares (OLS). We analyse DecoR theoretically and provide assumptions under which DecoR is consistent for estimating the causal effect.
Our approach was inspired by Mahecha et al. (2010) who predict temperature sensitivity of ecosystem respiratory processes in the case where basal respiration, the unobserved confounder, is slowly varying. While Mahecha et al. (2010) also consider the time series data in the frequency domain, they assume that the support of the confounder is known (which is not required in our framework). Furthermore, they focus on the application and neither provide a formal mathematical framework nor any statistical guarantees. The work by Ćevid et al. (2020) and Scheidegger et al. (2023) also consider sparsity with hidden confounders. However, they consider i.i.d. data and assume a sparse parameter vector rather than sparse confounding. Consequently, their procedure is different from ours in that they apply singular value decomposition to the data and then trim the large eigenvalues. A related line of research focuses on spatial deconfounding in environmental and epidemiological applications; see, e.g., Clayton et al. (1993); Paciorek (2010); Page et al. (2017). These works assume that a spatial regression problem is confounded by an unobserved confounder that is slowly varying. Different methods have been developed to solve this problem. In Reich et al. (2006), Hughes and Haran (2013) and Prates et al. (2019) the residual spatial process is restricted to be orthogonal to the covariates. Another approach is to remove the slowly varying components from either the response or the covariates or both (Paciorek, 2010; Thaden and Kneib, 2018; Keller and Szpiro, 2020; Dupont et al., 2022). A similar methodology, that also includes removing the slowly varying part of the covariate, was developed by Sippel et al. (2019) for meteorological time series data. More recently, Guan et al. (2023) and Marques et al. (2022) consider Gaussian random fields as a data generation process and propose to remove confounding by considering different scales when estimating the (unconfounded) covariance matrix.
@@
The remainder of this paper is structured as follows. We formally introduce the problem of causal inference with unobserved confounders for time series in Section 2. In Section 2.1 we present DecoR and showcase how it is used for causal inference in the presence of unobserved confounders. We provide theoretical guarantees for BFS and Torrent in Section 3 and guarantees for DecoR in Section 4. In Section 5 we provide experimental results on both synthetic and real-world data.
@@ 2 Sparse Representation for Deconfounding
Let $d\in\mathbb{N}$ and let $T\in\mathbb{R}_{>0}$ denote a fixed time horizon. Let $X=(X_{t})_{t\in[0,T]}$ be a stochastic process in $\mathbb{R}^{d}$ and $U=(U_{t})_{t\in[0,T]}$ a stochastic process in $\mathbb{R}$. Let $\eta=(\eta_{t})_{t\in[0,T]}$ be a process of i.i.d. centred Gaussian random variables independent of $X$ and with constant variance $\sigma^{2}_{\eta}\geq 0$. Let $\beta\in\mathbb{R}^{d}$ and let $Y=(Y_{t})_{t\in[0,T]}$ be a stochastic process that satisfies
@@
Consider a (known) orthonormal basis $\phi=\{\phi_{k}\}_{k\in\mathbb{N}}$ of $L^{2}([0,T])$. We typically choose the cosine basis (see Definition 2) as $\phi$ in our applications. The main assumption of this paper is that the confounder is sparse in this basis.
Definition 1 ($(\phi,G)$-sparse process ).
@@
The core idea is to consider the pairs $\{(T_{k}^{\phi,n}(X),T_{k}^{\phi,n}(Y))\}_{k\leq n}$ as a new data set and to observe that by Assumption 1 only the data points with an index $k\in G$ are confounded. If $G$ is known or if there exists a known subset $S\subseteq\mathbb{N}$ with $G\subseteq S$, the unconfounded data set $\{(T_{k}^{\phi,n}(X),T_{k}^{\phi,n}(Y))\}_{k\notin S}$ can be analysed using standard linear inference methods to, for example, consistently estimate $\beta$. However, in many cases, $G$ (or $S$) might be unknown. In the next section, we introduce DecoR, an algorithm for estimating $\beta$ without assuming knowledge of $G$ but assuming sparsity instead.
@@ 3 Theoretical Guarantees for Robust Regression
This work discusses two algorithms for estimating $\beta$ in (4), both of these methods include some form of linear regression. Algorithm 2 takes as input a collection $\mathcal{U}_{n}$ of candidate sets for the inliers $G_{n}^{\mathsf{c}}$. For example, we might have knowledge of an upper bound $a$ for the number of outliers. We can then define $\mathcal{U}_{n}=\{S\subseteq\{1,\dots,n\}\ |\ |S|=n-a\}$. We can then search over all elements in $\mathcal{U}_{n}$, compute the OLS estimate on the corresponding data, and choose the estimate with the smallest prediction error. We call this method Brute Force Search (BFS) and detail it in Algorithm 2. We will see in Section 3.1 that if, indeed, $G_{n}^{\mathsf{c}}\in\mathcal{U}_{n}$, the procedure is consistent under suitable assumptions.
@@
while $\left\lVert e_{S_{t}}\right\rVert_{2}<\mathrm{err}$ do
@@
     $v\leftarrow|Y-X\hat{\beta}^{t}_{\texttt{Tor}}|$
     $S_{t}\leftarrow\mathrm{HT}(v,a)$triangleright see (6)
     $e\leftarrow|Y_{S_{t}}-X_{S_{t}}\hat{\beta}^{t}_{\texttt{Tor}}|$
@@ 3.1 Guarantees for BFS
where $K>0$ is the constant from Lemma 2, and define
@@
The proof can be found in Appendix D.2. Theorem 1 implies that BFS is consistent if the noise variance $\sigma^{2}$ is equal to 00 and $\min_{S\in\mathcal{U}_{n}}||X_{S\setminus G_{n}}||_{2}\neq 0$, implying that $\mathcal{U}_{n}$ does not contain any subsets of $G_{n}$. Furthermore, it shows that if the fraction of outliers is vanishing with increasing $n$, BFS is consistent under mild assumptions on the collection $\mathcal{U}_{n}$ of candidate sets; it suffices, for example, if the distribution of $X$ is such that [^6] See Appendix 5 for a formal definition.,
$\min_{S\in\mathcal{U}_{n}}||X_{S\setminus G_{n}}||_{2}\in\Omega_{\mathbb{P}}(\sqrt{n}),\qquad\max_{S\in\mathcal{U}_{n}}||X_{S}||_{2}\in\mathcal{O}_{\mathbb{P}}(\sqrt{n})$ (7)
@@
For details, see Appendix E.2. Conditions (8) and (7) hold, for example, when we have a known sequence $\{c_{n}\}_{n\in\mathbb{N}}$ such that, for all $n$, $|G_{n}|\leq c_{n}$ and $c_{n}$ $\in o(\sqrt{n}/\log(n))$, we define the collection of candidate sets as $\mathcal{U}_{n}\coloneqq\{S\in\mathcal{P}(\{1,\dots,n\})\mid|S|=n-c_{n}\}$ and the rows of $X$ are i.i.d. Gaussian, see Bhatia et al. (2015, Theorem 15). If the fraction of outliers is non-vanishing (and $c_{n}<n/2$), Theorem 1 implies that for a distribution of $X$ satisfying (7) and the same choice of candidate sets defined by $\mathcal{U}_{n}\coloneqq\{S\in\mathcal{P}(\{1,\dots,n\})\mid|S|=n-c_{n}\}$, that $|\hat{\beta}^{n}_{\texttt{BFS}}(\mathcal{U}_{n})-\beta|\in\mathcal{O}(\sigma)$. Theorem 1 implies uniform consistency results, too. For example, in the case of $\sigma=0$, we obtain uniform consistency even over classes of distributions with outliers that shift the original data points by an arbitrary amount. On the contrary, regression that is based on a robust loss function, such as the Huber loss (Huber, 1964), does not come with similar guarantees, not even with respect to pointwise consistency or in the case $\sigma=0$.
@@ 3.2 Guarantees for Torrent
In practice, we observe that Torrent requires only very few iterations until it converges (cf. Table 9). We can use Lemma 1 to obtain an improved guarantee for Algorithm 3.
Theorem 2.
@@
$\eta\coloneqq\max_{S\subseteq\{1,\dots,n\}\text{ s.t. }|S|=a_{n}}\frac{\left\lVert X_{V(S)}\right\rVert_{2}}{\sqrt{\lambda_{\mathrm{min}}\left(X_{S}^{T}X_{S}\right)}}<\frac{1}{\sqrt{2}}.$ (9)
@@
$\left\lVert\hat{\beta}^{n,a_{n}}_{\texttt{Tor}}-\beta\right\rVert_{2}\leq\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\epsilon_{S_{t}}\right\rVert_{2}+\frac{\sqrt{2}\left\lVert X_{V(S_{t})}\right\rVert_{2}\left\lVert\left(X_{S_{t}}^{\top}X_{S_{t}}\right)^{-1}X_{S_{t}}^{\top}\epsilon_{S_{t}}\right\rVert_{2}+\sqrt{2}\|\epsilon_{V(S_{t})}\|_{2}}{\sqrt{\lambda_{\mathrm{min}}\left(X_{S_{t}}^{T}X_{S_{t}}\right)}\left(1-\sqrt{2}\eta\right)}.$
@@
Corollary 1 (sub-Gaussian noise).
Assume the setting of Theorem 2 (including the assumption that (9) holds), with i.i.d. zero-mean sub-Gaussian noise $\epsilon$ with variance proxy $\sigma^{2}$ and assume $c_{n}<n/2$. Then, there exists a constant $K>0$ such that for all $\delta>0$ with probability at least $1-\delta$
@@
While Corollary 1 states consistency under appropriate conditions and vanishing fractions of outliers, the results in Bhatia et al. (2015) are not sufficiently strong to imply consistency in this setting. Furthermore, even for non-vanishing fraction of outliers our bounds are in general tighter since the maximum over $S$ in the definition of $\eta$ is taken jointly over the denominator and numerator, not separately as in Bhatia et al. (2015).
@@ 4 Guarantees for DecoR
For the remainder of this section we assume Setting 1 and that Assumption 1 is satisfied for $G$ and $\phi$. More precisely, let $G\subseteq\mathbb{N}$, let $\phi=\{\phi_{k}\}_{k\in\mathbb{N}}$ be an orthonormal basis of $L^{2}([0,T])$, let $U=(U_{t})_{t\in[0,T]}\in\mathcal{M}^{2}_{[0,T]}$ be a $(\phi,G)$-sparse process and let $X=(X_{t})_{t\in[0,T]}$ be a stochastic process in $\mathbb{R}^{d}$. Let $\eta=(\eta_{j})_{j\in[0,T]}$ be a stochastic process of i.i.d. centred Gaussian random variables with variance $\sigma_{\eta}^{2}\geq 0$ and let $Y_{t}$ be a stochastic process, such that, for all $t\in[0,T]$,
@@
We also assume some regularity conditions on $\phi$ and $U$. Specifically, we assume that for all $k\in\mathbb{N}$ it holds that $\phi_{k}$ is right-continuous (or left-continuous) and that the trajectories of $U$ satisfy the same condition almost surely. This implies that almost surely888 Without assuming right-continuity (10) only holds almost surely, almost everywhere; because we need to discretize the process, we require equality almost surely for all $t\in[0,T]$. , for all $t\in[0,T]$
@@
    For all $n\in\mathbb{N}$ and $l,k\leq n$ we have $\frac{1}{n}\sum_{j=1}^{n}\phi_{l}(Tj/n)\phi_{k}(Tj/n)=\mathbbm{1}{\{l=k\}}$.
@@
    For every $\delta>0$ there exist $c^{\prime}>0$ and $\bar{n}\in\mathbb{N}$ such that for all $n\geq\bar{n}$ there exist $S^{\prime}_{n}\subseteq\{1,\dots,n\}$ with $|S^{\prime}_{n}|=2c_{n}+d$ such that for all $S^{\prime\prime}\subset S^{\prime}_{n}$ with $|S^{\prime\prime}|=d$ it holds with probability at least $1-\delta$
@@
Assumption 2 (i) states that the chosen orthonormal basis maintains its orthogonality and normalization properties when applied to discretized observations; this is satisfied, for example, by the cosine basis or the Haar basis (Haar, 1909), see Appendix C. Together with (10), this implies (2). Assumption 2 (ii) requires that $X$ is non-sparse in the $\phi$-domain. This condition intuitively says that there is more information in $X$ than is lost by confounding. This condition is relatively mild and accommodates a wide array of commonly used stochastic processes, including Ornstein-Uhlenbeck processes, Brownian motions, and band-limited processes. We are now able to state the main theoretical results, the convergence properties of DecoR-BFS (DecoR with BFS as the robust regression algorithm) and DecoR-Tor (DecoR with Tor as the robust regression algorithm).
Theorem 3 (Convergence properties of DecoR-BFS).
Let $c_{n}$ be a known sequence of natural numbers such that $|G_{n}|\leq c_{n}$ and let Assumption 2 be satisfied. Assume that $d=1$, that $U$ is independent of $\epsilon$, that $\sup_{t\in[0,T]}\mathbb{E}[X_{t}^{2}]<\infty$ and that for all $n\in\mathbb{N}$ and $j,i\leq T$ it holds that $\frac{1}{n}\sum_{l=1}^{n}\phi_{l}(Tj/n)\phi_{l}(Ti/n)=\mathbbm{1}{\{i=j\}}$. Define a sequence of candidate sets by $\mathcal{U}_{n}:=\{S\subseteq\{1,\dots,n\}\ |\ |S|=n-c_{n}\}$. If DecoR is executed with BFS and the sequence of sample sets $\mathcal{U}_{n}$, then
$\left|\hat{\beta}^{\phi,n}_{\texttt{DecoR}}-\beta\right|\in\mathcal{O}_{\mathbb{P}}\left(\sigma_{\eta}\left(\frac{c_{n}\log(n/c_{n})}{n}\right)^{1/4}\right).$
Many commonly used bases, such as the cosine basis and the Haar basis, satisfy the conditions of Theorem 3.
Theorem 4 (Convergence properties of DecoR-Tor).
@@
If DecoR is executed with Torrent and the sequence of threshold parameters $a_{n}=n-c_{n}$, then
$\left\lVert\hat{\beta}^{\phi,n}_{\texttt{DecoR}}-\beta\right\rVert_{2}$ $\in\mathcal{O}_{\mathbb{P}}\left(\sigma_{\eta}\sqrt{\frac{c_{n}\log(n/c_{n})}{n}}\right).$
If the assumptions of Theorem 3 or Theorem 4 are satisfied and either there is no noise, that is, $\sigma_{\eta}=0$, or the number of confounded components $c_{n}$ grows more slowly than the sample size $n$, that is, $c_{n}\in o(n)$, then DecoR produces a consistent estimator for $\beta$. Conversely, if there is non-zero noise and the number of confounded components $c_{n}$ is asymptotically proportional to $n$, that is, $c_{n}\sim n$, the estimation error of DecoR is asymptotically bounded by the noise variance $\sigma_{\eta}$, up to constant factors. While DecoR-Tor has polynomial-time complexity and in our experiments converges very fast (cf. Table 9), the theoretical result (Theorem 4) contains stronger theoretical assumptions than the one based on BFS (Theorem 3) , specifically it requires (11) to hold. In contrast, DecoR-BFS does not require (11), but its computational complexity is exponential in $n$ when $c_{n}$ is not bounded by a constant.